Suppose you deposit $2000 into an account at the end of each year for the next 10 years and then you stop making deposits. If the account earns 12% (assume annual compounding), how much will be in the account at the end of 30 years?

35,097

192,926

338,560

482,665

Insufficient information to compute
Answer #1

Value at year 10 = Annuity * [(1 + r)n - 1] / r

Value at year 10 = 2000 * [(1 + 0.12)10 - 1] / 0.12

Value at year 10 = 2000 * 17.548735

Value at year 10 = 35,097.47014

Value at year 30 = Present value (1 + r)n

Value at year 30 = 35,097.47014 (1 + 0.12)20

Value at year 30 = 35,097.47014 * 9.646293

Value at year 30 = 338,560